# 0-supple.tex

### v1

@@ C.2 Proof of the Results for the Heterogeneous Setting
Recall that $\Lambda_{j,\ell}$ denotes the singular value matrix of $\mathcal{M}_{j}(\mathcal{G}_{\ell})$. Let $\Lambda_{j,U}$ denote the top-left $r_{j,U}\times r_{j,U}$ diagonal block of $\Lambda_{j,\ell}$ and $\Lambda_{j,V_{\ell}}$ be the remaining block, i.e., $\Lambda_{j,\ell}=\begin{pmatrix}\Lambda_{j,U}&0\\ 0&\Lambda_{j,V_{\ell}}\end{pmatrix}$. By (8) and (20), the columns of $\widehat{U}_{j,\ell}$ are the first $r_{j,U}$ eigenvectors of
@@
$\sup_{j,\ell}\left\lVert\mathfrak{E}_{j,\ell}\right\rVert_{2}\leq C_{2}\lambda_{\min}\kappa_{0}\sigma\sqrt{p},$ (35)
$\sup_{j,\ell}\left\lVert\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}+\zeta^{\top}_{j,\ell,1}\right\rVert_{2}\leq C_{2}\kappa_{0}^{2}p\sigma^{2}\sqrt{r},$ (36)
@@
$\zeta_{j,\ell,1}=T^{*}_{j,\ell}[UV]_{-j,\ell}[UV]_{-j,\ell}^{\top}Z^{\top}_{j,\ell},$
and some absolute constant $C_{2}$, with probability tending to one.
@@
$\sup_{j,\ell}\left\lVert\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\sup_{j,\ell}\frac{\sqrt{r_{j,U}}\left\lVert\mathfrak{E}_{j,\ell}\right\rVert_{2}}{\lambda^{2}_{r_{j,U},j,\ell}-\lambda^{2}_{r_{j,U}+1,j,\ell}}\lesssim\frac{\sqrt{pr_{U}}\kappa\sigma}{\Delta},$ (37)
where $r_{U}=\max_{j}r_{j,U}$.
Define
$\Sigma_{j}^{*}=\frac{1}{L}\sum_{\ell=1}^{L}\mathbb{E}\left[\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}\right],$
and let $U_{j}^{*}=\mathrm{svd}_{r_{j,U}}(\Sigma_{j}^{*})$. By (37), we have
$\sup_{j,\ell}\left\lVert\mathbb{E}\left[\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}\right]-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\frac{\sqrt{pr_{U}}\kappa\sigma}{\Delta},$ (38)
and thus
$\left\lVert\Sigma_{j}^{*}-U_{j}U_{j}^{\top}\right\rVert_{2}\leq\left\lVert\Sigma_{j}^{*}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\leq\frac{\sqrt{pr_{U}}\kappa\sigma}{\Delta}.$
Under the assumption that $(\Delta/\sigma)\gtrsim\kappa\sqrt{pr}$, by Weyl’s inequality, we obtain that $\lambda_{r_{j,U}}(\Sigma_{j}^{*})-\lambda_{r_{j,U}+1}(\Sigma_{j}^{*})=1+o(1)$, and hence, by Davis-Kahan Theorem,
$\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}^{*}U_{j}^{*\top}\right\rVert_{\rm F}\lesssim\left\lVert\frac{1}{L}\sum_{\ell=1}^{L}\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}-\Sigma_{j}^{*}\right\rVert_{\rm F}\leq\frac{1}{L}\left\lVert\sum_{\ell=1}^{L}(1-\mathbb{E})\left[\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}\right]\right\rVert_{\rm F}.$
By the proof of Lemma 4 in Fan et al. (2019), we have that
$\left\lVert\sum_{\ell=1}^{L}(1-\mathbb{E})\left[\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}\right]\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\sqrt{L\sup_{\ell}\left\lVert(1-\mathbb{E})\left[\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}\right]\right\rVert_{\rm F}^{2}}\right)=O_{\mathbb{P}}\left(\frac{\sqrt{Lpr_{U}}\kappa\sigma}{\Delta}\right),$
where the second equality follows from (37) and (38). Therefore, we obtain that
$\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}^{*}U_{j}^{*\top}\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\frac{\sqrt{pr_{U}}\kappa\sigma}{\sqrt{L}\Delta}\right).$ (39)
Now we focus on bounding $\left\lVert U_{j}^{*}U_{j}^{*\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}$. Let the $(\bm{u}_{1,j,\ell},\dots,\bm{u}_{p_{j},j,\ell})$ be the eigenvectors of $U_{j}\Lambda_{j,U}^{2}U_{j}^{\top}+V_{j,\ell}\Lambda_{j,V_{\ell}}^{2}V_{j,\ell}^{\top}$ and the corresponding eigenvalues be $\lambda_{1,j,\ell}^{2}>\dots>\lambda_{p_{j},j,\ell}^{2}$. For $k=1,\dots,r_{j,U}$, let $O_{k,j,\ell}=\sum_{k^{\prime}=r_{j,U}+1}^{+\infty}(\lambda^{2}_{k^{\prime},j,\ell}-\lambda^{2}_{k,j,\ell})^{-1}\bm{u}_{k^{\prime}}\bm{u}_{k^{\prime}}^{\top}$, and define
$f_{j,\ell}:\mathbb{R}^{p_{j}\times r_{j,U}}\to\mathbb{R}^{p_{j}\times r_{j,U}},(\bm{w}_{1},\dots,\bm{w}_{r_{j,U}})\to(-O_{1,j,\ell}\bm{w}_{1},\dots,-O_{r_{j,U},j,\ell}\bm{w}_{r_{j,U}}).$
Since
$\sup_{j,\ell}\left\lVert\mathfrak{E}_{j,\ell}\right\rVert_{2}\lesssim\lambda_{\max}\sigma\sqrt{p}\lesssim\Delta^{2}\leq\lambda_{r_{j,U},j,\ell}^{2}-\lambda_{r_{j,U}+1,j,\ell}^{2},$
(using $(\Delta/\sigma)\gtrsim\kappa\sqrt{pr_{j,U}}$), we can apply Lemma 2 in Fan et al. (2019) and obtain that
$\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}=U_{j}U_{j}^{\top}+f_{j,\ell}(\mathfrak{E}_{j,\ell}U_{j})U_{j}^{\top}+U_{j}f_{j,\ell}(\mathfrak{E}_{j,\ell}U_{j})^{\top}+\mathfrak{R}_{j,\ell},$
where the remainder term satisfies
$\sup_{j,\ell}\left\lVert\mathfrak{R}_{j,\ell}\right\rVert_{\rm F}\lesssim\frac{\sqrt{r_{j,U}}\left\lVert\mathfrak{E}_{j,\ell}\right\rVert_{2}^{2}}{\Delta^{4}}\lesssim\frac{\sqrt{r_{U}}p\kappa^{2}\sigma^{2}}{\Delta^{2}}.$
By (36) and $\mathbb{E}[\zeta_{j,\ell,1}]=0$, we have that $\left\lVert\mathbb{E}\left[\mathfrak{E}_{j,\ell}\right]\right\rVert_{2}\lesssim\kappa_{0}^{2}p\sigma^{2}\sqrt{r}$, and hence
$\left\lVert\mathbb{E}\left[f_{j,\ell}(\mathfrak{E}_{j,\ell}U_{j})U_{j}^{\top}\right]\right\rVert_{\rm F}=\left\lVert f_{j,\ell}(\mathbb{E}[\mathfrak{E}_{j,\ell}U_{j}])\right\rVert_{\rm F}\leq\Delta^{-2}\left\lVert\mathbb{E}[\mathfrak{E}_{j,\ell}]U_{j}\right\rVert_{\rm F}\leq\Delta^{-2}\sqrt{r_{j,U}}\left\lVert\mathbb{E}[\mathfrak{E}_{j,\ell}]\right\rVert_{2}\lesssim\kappa_{0}^{2}pr\sigma^{2}\Delta^{-2}.$
Therefore, we obtain
$\left\lVert\mathbb{E}\left[\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}\right]-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\frac{pr(\kappa^{2}+\kappa_{0}^{2})\sigma^{2}}{\Delta^{2}},$
which implies that
$\left\lVert U_{j}^{*}U_{j}^{*\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\frac{pr(\kappa^{2}+\kappa_{0}^{2})\sigma^{2}}{\Delta^{2}}.$
Together with (39), we conclude that
$\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}=O_{\mathbb{P}}\left(\sqrt{\frac{pr}{L}}\frac{\kappa\sigma}{\Delta}+\frac{pr(\kappa^{2}+\kappa_{0}^{2})\sigma^{2}}{\Delta^{2}}\right).$
@@
Since
$\sup_{j,\ell}\left\lVert V_{j,\ell}\Lambda_{j,V_{\ell}}^{2}V_{j,\ell}^{\top}\left(U_{j}U_{j}^{\top}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\right)\right\rVert_{2}=O_{\mathbb{P}}\left(\sqrt{\frac{pr}{L}}\frac{\lambda^{2}_{\max}\kappa\sigma}{\Delta}+\frac{pr\lambda_{\max}^{2}(\kappa^{2}+\kappa_{0}^{2})\sigma^{2}}{\Delta^{2}}\right),$
@@
$=O_{\mathbb{P}}\left(\frac{pr}{L}\frac{\lambda^{2}_{\max}\kappa^{2}\sigma^{2}}{\Delta^{2}}+\frac{p^{2}r^{2}\lambda_{\max}^{2}(\kappa^{2}+\kappa_{0}^{2})^{2}\sigma^{4}}{\Delta^{4}}\right)$
$=O_{\mathbb{P}}\left(\sqrt{\frac{pr}{L}}\frac{\lambda^{2}_{\max}\kappa\sigma}{\Delta}+\frac{pr\lambda_{\max}^{2}(\kappa^{2}+\kappa_{0}^{2})\sigma^{2}}{\Delta^{2}}\right),$
@@
$\sup_{j,\ell}\left\lVert\left(I_{p_{j}}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\right)\mathfrak{E}_{j,\ell}\left(I_{p_{j}}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\right)\right\rVert_{2}=O_{\mathbb{P}}\left(\sqrt{p}\lambda_{\max}\sigma\right).$

### v3

@@ D.2 Proof of the Results for the Heterogeneous Setting
Recall that $\Lambda_{j,\ell}$ denotes the singular value matrix of $\mathcal{M}_{j}(\mathcal{G}_{\ell})=:G_{j,\ell}$. By (8), $G_{j,\ell}G_{j,\ell}^{\top}=\Lambda_{j,\ell}^{2}$. Let $\Lambda_{j,U}$ denote the top-left $r_{j,U}\times r_{j,U}$ diagonal block of $\Lambda_{j,\ell}$ and $\Lambda_{j,V_{\ell}}$ be the remaining block, i.e., $\Lambda_{j,\ell}=\begin{pmatrix}\Lambda_{j,U}&0\\ 0&\Lambda_{j,V_{\ell}}\end{pmatrix}$. By (8) and (25), the columns of $\widehat{U}_{j,\ell}$ are the first $r_{j,U}$ eigenvectors of
@@
$\sup_{j,\ell}\left\lVert\mathfrak{E}_{j,\ell}\right\rVert_{2}\leq C^{\prime}_{2}\lambda_{\min}\kappa_{0}\sigma\sqrt{p},$ (49)
$\sup_{j,\ell}\left\lVert\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}+\zeta^{\top}_{j,\ell,1}\right\rVert_{2}\leq C^{\prime}_{2}\kappa_{0}^{2}p\sigma^{2}\sqrt{r},$ (50)
@@
$\zeta_{j,\ell,1}=T^{*}_{j,\ell}[U\,V]_{-j,\ell}[U\,V]_{-j,\ell}^{\top}Z^{\top}_{j,\ell}=[U_{j}\,V_{j,\ell}]G_{j,\ell}[U\,V]_{-j,\ell}^{\top}Z_{j,\ell}^{\top},$
and some absolute constant $C^{\prime}_{2}$, with probability at least $1-C_{1}^{\prime}Le^{-c_{1}^{\prime}p}$.
Let the $(\bm{u}_{1,j,\ell},\dots,\bm{u}_{p_{j},j,\ell})$ be the eigenvectors of $U_{j}\Lambda_{j,U}^{2}U_{j}^{\top}+V_{j,\ell}\Lambda_{j,V_{\ell}}^{2}V_{j,\ell}^{\top}$ and the corresponding eigenvalues be $\lambda_{1,j,\ell}^{2}>\dots>\lambda_{p_{j},j,\ell}^{2}$. For $k=1,\dots,r_{j,U}$, let $O_{k,j,\ell}=\sum_{k^{\prime}=r_{j,U}+1}^{p_{j}}(\lambda^{2}_{k,j,\ell}-\lambda^{2}_{k^{\prime},j,\ell})^{-1}\bm{u}_{k^{\prime}}\bm{u}_{k^{\prime}}^{\top}$, and define
$f_{j,\ell}:\mathbb{R}^{p_{j}\times r_{j,U}}\to\mathbb{R}^{p_{j}\times r_{j,U}},(\bm{w}_{1},\dots,\bm{w}_{r_{j,U}})\to(O_{1,j,\ell}\bm{w}_{1},\dots,O_{r_{j,U},j,\ell}\bm{w}_{r_{j,U}}).$
Since, with probability at least $1-C_{1}^{\prime}Le^{-c_{1}^{\prime}p}$,
$\sup_{j,\ell}\left\lVert\mathfrak{E}_{j,\ell}\right\rVert_{2}\lesssim\kappa_{0}\lambda_{\min}\sigma\sqrt{p}\ll\lambda_{\min}\Delta\leq\lambda_{r_{j,U},j,\ell}^{2}-\lambda_{r_{j,U}+1,j,\ell}^{2},$
(using $(\Delta/\sigma)\gg\sqrt{p}$ and $\kappa_{0}=O(1)$), we can apply Lemma 2 in Fan et al. (2019) and obtain that
$\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}^{\top}=U_{j}U_{j}^{\top}+f_{j,\ell}(\mathfrak{E}_{j,\ell}U_{j})U_{j}^{\top}+U_{j}f_{j,\ell}(\mathfrak{E}_{j,\ell}U_{j})^{\top}+\mathfrak{R}_{j,\ell},$ (51)
where the remainder term satisfies
$\sup_{j,\ell}\left\lVert\mathfrak{R}_{j,\ell}\right\rVert_{\rm F}\lesssim\frac{\sqrt{r_{j,U}}\left\lVert\mathfrak{E}_{j,\ell}\right\rVert_{2}^{2}}{\Delta^{2}\lambda_{\min}^{2}}\lesssim\frac{\sqrt{r}p\kappa_{0}^{2}\sigma^{2}}{\Delta^{2}}.$ (52)
By the linearity of $f_{j,\ell}$, we further have
$f_{j,\ell}(\mathfrak{E}_{j,\ell}U_{j})=f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}+f_{j,\ell}\big{[}\zeta_{j,\ell,1}^{\top}U_{j}\big{]}+f_{j,\ell}\big{[}(\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top})U_{j}\big{]}.$
Note that $\left\lVert f_{j,\ell}(X)\right\rVert_{\rm F}\leq\frac{1}{\inf_{j,\ell}(\lambda^{2}_{r_{j,U},j,\ell}-\lambda^{2}_{r_{j,U}+1,j,\ell})}\left\lVert X\right\rVert_{\rm F}$ and $\lambda^{2}_{r_{U},j,\ell}-\lambda^{2}_{r_{U}+1,j,\ell}\geq\Delta^{2}$. Hence
$\sup_{j,\ell}\left\lVert f_{j,\ell}\big{[}(\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top})U_{j}\big{]}\right\rVert_{\rm F}\leq\Delta^{-2}\sup_{j,\ell}\left\lVert(\mathfrak{E}_{j,\ell}-\zeta_{j,\ell,1}-\zeta_{j,\ell,1}^{\top})U_{j}\right\rVert_{\rm F}\lesssim\kappa_{0}^{2}pr\sigma^{2}\Delta^{-2}.$ (53)
Now we bound $\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}$ and $\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}^{\top}U_{j}\big{]}$. Since $\mathrm{vec}(G_{j,\ell}[U\,V]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j})=(U_{j}^{\top}\otimes G_{j,\ell}[U\,V]_{-j,\ell}^{\top})\mathrm{vec}(Z_{j,\ell}^{\top})$, we have that
$\mathrm{vec}(G_{j,\ell}[U\,V]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j})\sim\mathcal{N}(0,\sigma^{2}(I_{r_{j,U}}\otimes\Lambda_{j,\ell}^{2})),$
which shows that $G_{j,\ell}[U\,V]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j}$ has independent normal entries with variance at most $\lambda_{\max}^{2}\sigma^{2}$. Let $G_{j,\ell}[U\,V]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j}=(\bm{w}_{1,j,\ell},\dots,\bm{w}_{r_{j,U},j,\ell})$. Since
$\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}=\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}([U_{j}\,V_{j,\ell}]G_{j,\ell}[U\,V]_{-j,\ell}^{\top}Z_{j,\ell}^{\top}U_{j}),$
we can write
$\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}=\left(\frac{1}{L}\sum_{\ell=1}^{L}O_{1,j,\ell}[U_{j}\,V_{j,\ell}]\bm{w}_{1,j,\ell},\dots,\frac{1}{L}\sum_{\ell=1}^{L}O_{r_{j,U},j,\ell}[U_{j}\,V_{j,\ell}]\bm{w}_{r_{j,U},j,\ell}\right).$
In the above matrix, each column $\frac{1}{L}\sum_{\ell=1}^{L}O_{k,j,\ell}[U_{j}V_{j,\ell}]\bm{w}_{k,j,\ell}$ is independent among different $k$ and has norm distribution $\mathcal{N}\left(0,\Sigma_{j,k}\right)$, where
$\Sigma_{j,k}=\frac{\sigma^{2}}{L^{2}}\sum_{\ell=1}^{L}O_{1,j,\ell}[U_{j}V_{j,\ell}]\Lambda_{j,\ell}^{2}[U_{j}V_{j,\ell}]^{\top}O_{1,j,\ell}^{\top}.$
Therefore, $\mathrm{vec}\left(\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right)\sim\mathcal{N}(0,\Sigma_{j})$, where $\Sigma_{j}=\mathrm{diag}(\Sigma_{j,1},\Sigma_{j,2},\dots,\Sigma_{j,r_{j,U}})$. Since
$\left\lVert\Sigma_{j}\right\rVert_{2}\leq\max_{k}\left\lVert\Sigma_{j,k}\right\rVert_{2}\leq L^{-1}\sigma^{2}\lambda^{2}_{\max}\lambda_{\min}^{-2}\Delta^{-2},$
we obtain that
$\left\lVert\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}U_{j}^{\top}\right\rVert_{\rm F}$ $=\left\lVert\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right\rVert_{\rm F}$ (54)
$=\left\lVert\mathrm{vec}\left(\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right)\right\rVert_{2}$
$\leq\left\lVert\Sigma_{j}^{1/2}\right\rVert_{2}\left\lVert\Sigma_{j}^{-1/2}\mathrm{vec}\left(\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right)\right\rVert_{2}$
$\leq\sqrt{2pr/L}\kappa_{0}\sigma\Delta^{-1},$
with probability at least $1-2e^{-pr/8}$, where we apply (41) to bound the norm of the vector $\Sigma_{j}^{-1/2}\mathrm{vec}\left(\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}U_{j}\big{]}\right)\sim\mathcal{N}(0,I_{p_{j}r_{j,U}})$.
A similar argument can be used to bound
$\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}^{\top}U_{j}\big{]}=\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}(Z_{j,\ell}[UV]_{-j,\ell}G_{j,\ell}^{\top}[U_{j}V_{j,\ell}]^{\top}U_{j}),$
since $Z_{j,\ell}[UV]_{-j,\ell}G_{j,\ell}^{\top}[U_{j}V_{j,\ell}]^{\top}U_{j}=Z_{j,\ell}[UV]_{-j,\ell}G_{j,\ell}^{\top}\begin{pmatrix}I_{r_{j,U}}\\ \bm{0}_{r_{j,U},r_{j,V,\ell}}\end{pmatrix}$, and $Z_{j,\ell}[UV]_{-j,\ell}G_{j,\ell}^{\top}$ has independent normal entries with variance at most $\lambda_{\max}^{2}\sigma^{2}$. Then repeating the previous procedure leads to
$\left\lVert\frac{1}{L}\sum_{\ell=1}^{L}f_{j,\ell}\big{[}\zeta_{j,\ell,1}^{\top}U_{j}\big{]}U_{j}^{\top}\right\rVert_{\rm F}\leq\sqrt{2pr/L}\kappa_{0}\sigma\Delta^{-1},$ (55)
with probability at least $1-2e^{-pr/8}$.
Combing (51), (52), (53), (54), and (55), we obtain that
$\left\lVert\frac{1}{L}\sum_{\ell=1}^{L}\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\lesssim\sqrt{\frac{pr}{L}}\frac{\kappa_{0}\sigma}{\Delta}+\frac{\kappa_{0}^{2}pr\sigma^{2}}{\Delta^{2}},$
with probability at least $1-C_{1}^{\prime\prime}Le^{c_{1}^{\prime\prime}p}$ for some constants $c_{1}^{\prime\prime}$ and $C_{1}^{\prime\prime}$.
@@
$\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\leq 2\left\lVert\frac{1}{L}\sum_{\ell=1}^{L}\widehat{U}_{j,\ell}\widehat{U}_{j,\ell}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}.$
Therefore, we conclude that
$\left\lVert\widehat{U}_{j}\widehat{U}_{j}^{\top}-U_{j}U_{j}^{\top}\right\rVert_{\rm F}\leq\widetilde{C}_{2}\left(\sqrt{\frac{pr}{L}}\frac{\kappa_{0}\sigma}{\Delta}+\frac{pr\kappa_{0}^{2}\sigma^{2}}{\Delta^{2}}\right),$
with probability $1-C_{1}^{\prime\prime}Le^{-c_{1}^{\prime\prime}p}\geq 1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$, for some constants $\widetilde{C}_{1},\widetilde{c}_{1},\widetilde{C}_{2}$.
@@
Since, with probability at least $1-\widetilde{C}_{1}e^{-\widetilde{c}_{1}p}$,
$\sup_{j,\ell}\left\lVert V_{j,\ell}\Lambda_{j,V_{\ell}}^{2}V_{j,\ell}^{\top}\left(U_{j}U_{j}^{\top}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\right)\right\rVert_{2}\leq\widetilde{C}_{2}\lambda_{\max}^{2}\left(\sqrt{\frac{pr}{L}}\frac{\kappa_{0}\sigma}{\Delta}+\frac{pr\kappa_{0}^{2}\sigma^{2}}{\Delta^{2}}\right),$
@@
$\lesssim\frac{pr}{L}\frac{\lambda^{2}_{\max}\kappa_{0}^{2}\sigma^{2}}{\Delta^{2}}+\frac{p^{2}r^{2}\lambda_{\max}^{2}\kappa_{0}^{4}\sigma^{4}}{\Delta^{4}}$
$\ll\sqrt{\frac{pr}{L}}\frac{\lambda^{2}_{\max}\kappa_{0}\sigma}{\Delta}+\frac{pr\lambda_{\max}^{2}\kappa_{0}^{2}\sigma^{2}}{\Delta^{2}},$
@@
$\sup_{j,\ell}\left\lVert\left(I_{p_{j}}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\right)\mathfrak{E}_{j,\ell}\left(I_{p_{j}}-\widehat{U}_{j}\widehat{U}_{j}^{\top}\right)\right\rVert_{2}\lesssim\sqrt{p}\lambda_{\max}\sigma.$